Exercise 12-6 Pharoah Company, organized in 2016, has set up a single account for all intangible assets. The following summary discloses the debit entries that have been recorded during 2017. 1/2/17 4/1/17 7/1/17 8/1/17 9/1/17 Purchased patent (9-year life) Purchase goodwill (indefinite life) Purchased franchise with 10-year life; expiration date 7/1/27 Payment of copyright (5-year life) Research and development costs $475,200 365,600 456,000 120,000 241,000 $1,657,800 Prepare the necessary entry to clear the Intangible Assets account and to set up separate accounts for distinct types of intangibles. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter 0 for the amounts.) Account Titles and Explanation Debit Credit

Make the entry as of December 31, 2017, recording any necessary amortization. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter o for the amounts.) Account Titles and Explanation Debit Credit

Reflect all balances accurately as of December 31, 2017 Use straight-line amortization

This question has three parts. thank you.
Answer #1
1
Debit Credit
Patents 475200
Goodwill 365600
Franchise 456000
Copyright 120000
Research and Development Expenses 241000
          Intangible Assets 1657800
2
Debit Credit
Amortization Expense 85600
        Patents 52800 =475200/9
        Franchise 22800 =456000/10*6/12
        Copyright 10000 =120000/5*5/12
3
Use straight-line amortization:
Patents 422400 =475200-52800
Goodwill 365600
Franchise 433200 =456000-22800
Copyright 110000 =120000-10000